# {tocentry}[Uncaptioned image]

### v1

@@ 1 Introduction
Many-body perturbation theory $GW$ approximation remains the gold-standard for predicting quasiparticle (QP) energies in weakly correlated solids and molecules 1, 2, 3, 4, 5, 6, 7. By dressing the one-electron Green’s function $G$ with a dynamically screened Coulomb interaction $W$ 8, 9, 10, the method captures exchange and long-range correlation beyond density-functional theory, routinely achieving high accuracy for band gaps, ionization energies, and electron affinities 1, 11, 12, 13. However, the scaling of $GW$ method is steep 3, 6.
Stochastic GWGW (sGWGW) was introduced to overcome the steep cost of conventional $GW$ implementations by replacing explicit summations over empty states with random sampling 14, 15, 16, 17. In this framework, Green’s functions and $W$ are evaluated with stochastic orbitals, reducing the scaling to nearly linear in system size while retaining controllable statistical error bars.
Although some deterministic $GW$ calculations have been implemented 18, 19, 20, most sGWGW studies have focused on non-magnetic or collinearly spin-polarised systems, leaving large-scale materials with non-collinear spins or strong spin–orbit coupling beyond reach. Such systems—ranging from skyrmionic magnets to topological insulators—require a spinor treatment in which wavefunctions are complex two-component objects.
Here we develop and test the spin-resolved sGWGW that unifies collinear and fully non-collinear spin configurations within a single stochastic framework. For collinear magnets, we retain separate stochastic sampling of majority and minority channels. For non-collinear cases, we introduce a complex stochastic orbital basis that preserves the real-value stochastic external charge applied at time zero under time-dependent Hartree (TDH) propagation 21, 9. This basis enables an unbiased evaluation of the random-phase-approximation effective screened interaction for spinors. By incorporating collinear and fully non-collinear spin configurations into a single, scalable scheme, our approach makes many-body quasiparticle (QP) calculations practical for large magnetic and spin–orbit–rich materials.
@@ 2 Methodology
As the implementation is based on earlier works 14, 15, 16, 17, we will here review only the key novel aspects of the computational methodology. The key is the recasting of the operator actions as sampling via random operators. This is accomplished in real space by introducing a random sampling of real-space coordinates and spin channels, with a random sign at each grid point
@@ 2.1 Green’s function sampling
The sampling of the non-interacting Kohn-Sham Green’s function is the key component of the stochastic perturbative treatment (e.g., sGWGW14, 15, 16, 17. Instead of representing $G$ via its individual orbital (basis) components, we express it as a statistical average over random functions, here generalized to spinors. Formally,
@@
where $|\zeta(r\sigma,t)\rangle\equiv i\hat{G}(t)|\bar{\zeta}(r\sigma)\rangle$, in which we employ the split-operator (Trotter) time propagation method as in 22. This equation is central to the stochastic $GW$ method, as it rewrites the Green’s function as a sum of separable terms involving random functions.
@@
Here, $W$ is the screened Coulomb interaction. In the random-phase approximation (consistent with the $GW$ approximation 23, 24, the induced potential is described by the fluctuations of the total charge density (computed from the time-dependent Hartree calculations detailed below)21, 9. As such, $u(r\sigma^{\prime},t)$ affects both spin channels identically and depends only on the perturbing state $\phi_{\sigma^{\prime}}(r^{\prime})$.
@@ 2.2 Real-time evaluation of screened interactions
The screened Coulomb interaction is conventionally split between the (dominant) exchange term computed explicitly and the polarization part due to the induced charge fluctuations. We will detail the calculations of the latter. Here, the time-ordered potential, which enters Eq. 5, is obtained by a simple transformation (detailed in 16) of its retarded counterpart polarization potential:
@@
This term is computed from time evolution of the charge density as $W$ is obtained as the potential due to the charge density fluctuations induced by the addition/removal of the electron. The details are provided in  16.
@@
We can also check that this new basis satisfies $\{\bar{\zeta}^{\prime*}(\mathbf{r}\sigma)\bar{\zeta}^{\prime}(\mathbf{r}^{\prime}\sigma^{\prime})\}=(dV)^{-1}\delta_{\mathbf{r}}$. Then the new source charge is $\bar{\zeta}^{\prime}(\mathbf{r})\phi(\mathbf{r})=\bar{\zeta}(\mathbf{r})|\phi(\mathbf{r})|$. This greatly simplifies the approach and it is possible to employ the same (TDH) approach as for the spin-independent evolution used in the past. Note that the additional complex phase $\phi/|\phi|$ is trivially absorbed in the green function $G$. An alternative is to compute the real and imaginary parts of the charge response separately 25, but this is significantly more computationally demanding. We therefore implement and test the $\phi$-dependent stochastic-basis method for the remainder of the paper.
@@ 3 Convergence of the real-time stochastic formalism
Deterministic $GW$ calculations require explicit construction and manipulation of large Green’s function and screened interaction matrices, with computational cost that grows steeply ($\mathcal{O}(N^{4})$) with system size 3, 6. By contrast, the sGWGW formalism avoids this bottleneck: the cost scales nearly linearly with the number of electrons 14, 15, 16, since operator actions are estimated through random sampling rather than full matrix operations. As we show below, the advantages of random sampling prevail even when considering systems with broken spin degeneracies, and, in fact, the stochastic formalism has a further significant reduction in the computational cost compared to the conventional (deterministic) calculations.
To quantify the scaling of the spin-resolved stochastic scheme, we focus on the dominant computational step: the real-time propagation of the electronic states, which are practically sampled by random vectors. To analyze the code performance, we estimate the time-cost ratios for non-spin, spin-collinear, and spin-noncollinear sGWGW. We find that for the most time-consuming step, the linear scaling in the non-spin polarized sGWGW are preserve. Moreover, because of self-averaging in sGWGW, spin-resolved calculations require fewer random functions $N_{\eta}$ for sampling, leading to only a modest increase in time cost compared to non-spin-polarized sGWGW (Table 1). In contrast, the conventional $GW$ implementations (in the energy domain), the self-energy calculation normally takes four times longer due to double the number of bands and double the size of each wave function 20, 26. This is governed by the size of the operator matrices and the corresponding vectors. The spinor Hamiltonian $H^{\sigma_{1}\sigma_{2}}_{ij}$ and self energy $\Sigma(\omega)^{\sigma_{1}\sigma_{2}}_{ij}$ matrices are 4 times larger with spinor $\left|\phi^{\sigma}_{j}\right\rangle$ basis. Note that this is irrespective of the underlying physical interactions, i.e., if the calculation is performed in the enlarged basis, this leads to redundant computations which add to the total cost of the calculation.
@@
Where $N_{occ}$ is the number of occupied states. As the system size increases, their values are kept bounded at $\sqrt{2/N_{\eta}}$. So only a small constant $N_{\eta}$ is required to achieve a small relative time-dependent density and thus a converged self-energy calculation. It has been illustrated in Refs. 16, 17.
@@ 4 Simulation and discussion
The calculations were carried out on the Perlmutter supercomputer at NERSC. Each CPU node is equipped with two AMD Milan processors (64 cores total) and 512 GB of memory. The implementation is trivially parallelized over stochastic orbitals and time steps, allowing efficient scaling across multiple nodes as in Ref. 16. For molecules, the Martyna-Tuckerman approach is used to avoid periodic images 27. We apply Coulomb-interaction truncation along $z$ and 2D periodic boundary conditions for 2D systems 28, 29.
@@ 4.1 Collinear magnetic systems
For spin-collinear configurations, we verify the spin-resolved stochastic $GW$ method using the prototypical ferromagnetic monolayer CrI3. It is a van der Waals layered material that hosts long-range ferromagnetic order down to the monolayer limit, making it a cornerstone of two-dimensional magnetism research 30, 31. Its relatively simple crystal structure and Ising-like magnetic anisotropy provide a well-defined platform for testing electronic-structure methods. Our aim here is to demonstrate that the stochastic $GW$ framework accurately captures quasiparticle energies in such 2D ferromagnets, thereby establishing its applicability to spin-polarized layered systems.
@@ 4.2.1 Finite systems
We validated the noncollinear sGWGW implementation on two finite systems with strong spin–orbit coupling (SOC): the I2 molecule and the Pb4S4 nanocrystal that are previously used in deterministic $GW$ benchmarks 19. To check robustness against numerical choices, we employed dense real-space grids ($dx=0.27$ a.u. for I2; $dx=0.3$ a.u. for Pb4S4) and two different Perdew-Burke-Ernzerhof (PBE) functional scalar-relativistic (SR) and fully relativistic (FR) norm-conserving pseudopotentials, PseudoDojo 33 and SG15 33, 34, 19.
For I2, we computed SR and FR quasiparticle energies using deterministic TDH propagation to build the screening (Table 3). Relative to SR $G_{0}W_{0}$, the FR treatment—which includes SOC and noncollinear spinor dynamics—enhances screening and systematically lowers the ionization energy, in agreement with deterministic FR $GW$ results reported in Ref. 19.
We then examined the Pb4S4 nanocrystal, which preserves the rock-salt motif of bulk PbS. The calculations show noticeable differences in the absolute QP energies that can be attributed to the use of different pseudopotential versions, which may introduce variations of up to a few hundred meV 19. We also see the difference between using PseudoDojo and SG15 pseudopotentials in Table 3 and 4. Further differences come from distinct real-space grid treatments and employing the real-time (time-domain) $GW$ builds screening via TDH propagation, in contrast to (truncated) summation over states methods 35. Nevertheless, we can reliably confirm the role of the relativistic effects. The FR–SR correction to the HOMO–LUMO gap is $\Delta E_{g}$ around 0.87 to 0.90 eV regardless of the pseudopotential used (Table 4), and it is consistent with deterministic $G_{0}W_{0}$ benchmarks (0.88 eV) 19.
@@ 4.2.2 Periodic system
For the periodic system, we verify the spin-resolved sGWGW method on bulk AlSb. A dense real-space grid with spacing $dx=0.27$ a.u. is employed, using PseudoDojo 33 PBE SR and FR pseudopotentials. The band gap calculations are shown in Table 5 and are consistent with deterministic $GW$ results 19.
We examine the convergence with respect to $N_{\eta}$, as shown in Fig.5. Both SR and FR pseudopotentials exhibit similar convergence, and $N_{\eta}=8$ is sufficient to obtain an accurate CBM quasiparticle energy $E_{qp}$ (Fig.2, top). This confirms the stochastic TDH performance analysis: noncollinear nonmagnetic systems require a comparable $N_{\eta}$ to achieve the same error level as collinear systems. The corresponding total core-hour costs are presented in Fig.2, bottom. In the noncollinear case, the computational effort increases for two reasons: (i) the FFTs must be performed on spinor wavefunctions, effectively doubling their cost, and (ii) the SOC nonlocal pseudopotential terms introduce additional overhead. As a result, the runtime is roughly three times that of the spin-unpolarized case, rather than just twice. Nonetheless, this is still significantly more efficient than deterministic $GW$, where achieving a similar accuracy would require a factor of four increase in additional cost. Further, note that this is a cost increase relative to the spin-degenerate calculation. In the sGWGW calculation, it scales merely linearly with a small prefactor, compared to the conventional calculations. Thus, the reported calculations are orders of magnitude less expensive than those obtained by a conventional route. This suggests that noncollinear sGWGW preserves the favorable linear-scaling behavior observed in spin-unpolarized systems, consistent with the analysis in Refs.14, 16.
@@ 5 Conclusion and outlook
In the future, this unified sGWGW framework will lay the foundation for systematically incorporating spin–spin interactions and vertex corrections. Such extensions will be crucial for accurately capturing excitonic effects, correlated magnetism, and unconventional superconductivity in real materials. Besides, AI-based solvers and learning modules offer a natural path to further accelerate stochastic propagations, particularly by exploiting low-rank structures and transferable features across materials 36, 37, 38, 39. These developments highlight the synergy between stochastic methods and data-driven approaches, opening the door to large-scale quasiparticle studies of large scale complex spin–orbit–coupled and magnetic materials that would otherwise be computationally prohibitive.

### v2

# Cost Reduction in Spin-dependent Stochastic $GW$ Calculations

@@ 1 Introduction
Many-body perturbation theory $GW$ approximation remains the gold-standard for predicting quasiparticle (QP) energies in weakly correlated solids and molecules 1, 2, 3, 4, 5, 6, 7, 8, 9, 10. By dressing the one-electron Green’s function $G$ with a dynamically screened Coulomb interaction $W$ 1, 11, 12, the method captures exchange and long-range correlation beyond density-functional theory, routinely achieving high accuracy for band gaps, ionization energies, and electron affinities 4, 13, 14, 15. However, the scaling of $GW$ method is steep 6, 9.
Stochastic GWGW (sGWGW) was introduced to overcome the steep cost of conventional $GW$ implementations by replacing explicit summations over empty states with random sampling 16, 17, 18, 19. In this framework, Green’s functions and $W$ are evaluated with stochastic orbitals, reducing the scaling to nearly linear in system size while retaining controllable statistical error bars.
Although some deterministic $GW$ calculations have been implemented 20, 21, 22, most sGWGW studies have focused on non-magnetic or collinearly spin-polarised systems, leaving large-scale materials with non-collinear spins or strong spin–orbit coupling beyond reach. Such systems—ranging from skyrmionic magnets to topological insulators—require a spinor treatment in which wavefunctions are complex two-component objects.
Here we develop and test the spin-resolved sGWGW that unifies collinear and fully non-collinear spin configurations within a single stochastic framework. For collinear magnets, we retain separate stochastic sampling of majority and minority channels. For non-collinear cases, we introduce a complex stochastic orbital basis that preserves the real-value stochastic external charge applied at time zero under time-dependent Hartree (TDH) propagation 23, 11. This basis enables an unbiased evaluation of the random-phase-approximation effective screened interaction for spinors. By incorporating collinear and fully non-collinear spin configurations into a single, scalable scheme, our approach makes many-body quasiparticle (QP) calculations practical for large magnetic and spin–orbit–rich materials.
@@ 2 Methodology
As the implementation is based on earlier works 16, 17, 18, 19, we will here review only the key novel aspects of the computational methodology. The key is the recasting of the operator actions as sampling via random operators. This is accomplished in real space by introducing a random sampling of real-space coordinates and spin channels, with a random sign at each grid point
@@ 2.1 Green’s function sampling
The sampling of the non-interacting Kohn-Sham Green’s function is the key component of the stochastic perturbative treatment (e.g., sGWGW16, 17, 18, 19. Instead of representing $G$ via its individual orbital (basis) components, we express it as a statistical average over random functions, here generalized to spinors. Formally,
@@
where $|\zeta(r\sigma,t)\rangle\equiv i\hat{G}(t)|\bar{\zeta}(r\sigma)\rangle$, in which we employ the split-operator (Trotter) time propagation method as in 24. This equation is central to the stochastic $GW$ method, as it rewrites the Green’s function as a sum of separable terms involving random functions.
@@
Here, $W$ is the screened Coulomb interaction. In the random-phase approximation (consistent with the $GW$ approximation 25, 26, the induced potential is described by the fluctuations of the total charge density (computed from the time-dependent Hartree calculations detailed below)23, 11. As such, $u(r\sigma^{\prime},t)$ affects both spin channels identically and depends only on the perturbing state $\phi_{\sigma^{\prime}}(r^{\prime})$.
@@ 2.2 Real-time evaluation of screened interactions
The screened Coulomb interaction is conventionally split between the (dominant) exchange term computed explicitly and the polarization part due to the induced charge fluctuations. We will detail the calculations of the latter. Here, the time-ordered potential, which enters Eq. 5, is obtained by a simple transformation (detailed in 18) of its retarded counterpart polarization potential:
@@
This term is computed from time evolution of the charge density as $W$ is obtained as the potential due to the charge density fluctuations induced by the addition/removal of the electron. The details are provided in  18.
@@
We can also check that this new basis satisfies $\{\bar{\zeta}^{\prime*}(\mathbf{r}\sigma)\bar{\zeta}^{\prime}(\mathbf{r}^{\prime}\sigma^{\prime})\}=(dV)^{-1}\delta_{\mathbf{r}}$. Then the new source charge is $\bar{\zeta}^{\prime}(\mathbf{r})\phi(\mathbf{r})=\bar{\zeta}(\mathbf{r})|\phi(\mathbf{r})|$. This greatly simplifies the approach and it is possible to employ the same (TDH) approach as for the spin-independent evolution used in the past. Note that the additional complex phase $\phi/|\phi|$ is trivially absorbed in the green function $G$. An alternative is to compute the real and imaginary parts of the charge response separately 27, but this is significantly more computationally demanding. We therefore implement and test the $\phi$-dependent stochastic-basis method for the remainder of the paper.
@@ 3 Convergence of the real-time stochastic formalism
Deterministic $GW$ calculations require explicit construction and manipulation of large Green’s function and screened interaction matrices, with computational cost that grows steeply ($\mathcal{O}(N^{4})$) with system size 6, 9. By contrast, the sGWGW formalism avoids this bottleneck: the cost scales nearly linearly with the number of electrons 16, 17, 18, since operator actions are estimated through random sampling rather than full matrix operations. As we show below, the advantages of random sampling prevail even when considering systems with broken spin degeneracies, and, in fact, the stochastic formalism has a further significant reduction in the computational cost compared to the conventional (deterministic) calculations.
To quantify the scaling of the spin-resolved stochastic scheme, we focus on the dominant computational step: the real-time propagation of the electronic states, which are practically sampled by random vectors. To analyze the code performance, we estimate the time-cost ratios for non-spin, spin-collinear, and spin-noncollinear sGWGW. We find that for the most time-consuming step, the linear scaling in the non-spin polarized sGWGW are preserve. Moreover, because of self-averaging in sGWGW, spin-resolved calculations require fewer random functions $N_{\eta}$ for sampling, leading to only a modest increase in time cost compared to non-spin-polarized sGWGW (Table 1). In contrast, the conventional $GW$ implementations (in the energy domain), the self-energy calculation normally takes four times longer due to double the number of bands and double the size of each wave function 22, 28. This is governed by the size of the operator matrices and the corresponding vectors. The spinor Hamiltonian $H^{\sigma_{1}\sigma_{2}}_{ij}$ and self energy $\Sigma(\omega)^{\sigma_{1}\sigma_{2}}_{ij}$ matrices are 4 times larger with spinor $\left|\phi^{\sigma}_{j}\right\rangle$ basis. Note that this is irrespective of the underlying physical interactions, i.e., if the calculation is performed in the enlarged basis, this leads to redundant computations which add to the total cost of the calculation.
@@
Where $N_{occ}$ is the number of occupied states. As the system size increases, their values are kept bounded at $\sqrt{2/N_{\eta}}$. So only a small constant $N_{\eta}$ is required to achieve a small relative time-dependent density and thus a converged self-energy calculation. It has been illustrated in Refs. 18, 19.
@@ 4 Simulation and discussion
The calculations were carried out on the Perlmutter supercomputer at NERSC. Each CPU node is equipped with two AMD Milan processors (64 cores total) and 512 GB of memory. The implementation is trivially parallelized over stochastic orbitals and time steps, allowing efficient scaling across multiple nodes as in Ref. 18. For molecules, the Martyna-Tuckerman approach is used to avoid periodic images 29. We apply Coulomb-interaction truncation along $z$ and 2D periodic boundary conditions for 2D systems 30, 31.
@@ 4.1 Collinear magnetic systems
For spin-collinear configurations, we verify the spin-resolved stochastic $GW$ method using the prototypical ferromagnetic monolayer CrI3. It is a van der Waals layered material that hosts long-range ferromagnetic order down to the monolayer limit, making it a cornerstone of two-dimensional magnetism research 32, 33. Its relatively simple crystal structure and Ising-like magnetic anisotropy provide a well-defined platform for testing electronic-structure methods. Our aim here is to demonstrate that the stochastic $GW$ framework accurately captures quasiparticle energies in such 2D ferromagnets, thereby establishing its applicability to spin-polarized layered systems.
@@ 4.2.1 Finite systems
We validated the noncollinear sGWGW implementation on two finite systems with strong spin–orbit coupling (SOC): the I2 molecule and the Pb4S4 nanocrystal that are previously used in deterministic $GW$ benchmarks 21. To check robustness against numerical choices, we employed dense real-space grids ($dx=0.27$ a.u. for I2; $dx=0.3$ a.u. for Pb4S4) and two different Perdew-Burke-Ernzerhof (PBE) functional scalar-relativistic (SR) and fully relativistic (FR) norm-conserving pseudopotentials, PseudoDojo 35 and SG15 35, 36, 21.
For I2, we computed SR and FR quasiparticle energies using deterministic TDH propagation to build the screening (Table 3). Relative to SR $G_{0}W_{0}$, the FR treatment—which includes SOC and noncollinear spinor dynamics—enhances screening and systematically lowers the ionization energy, in agreement with deterministic FR $GW$ results reported in Ref. 21.
We then examined the Pb4S4 nanocrystal, which preserves the rock-salt motif of bulk PbS. The calculations show noticeable differences in the absolute QP energies that can be attributed to the use of different pseudopotential versions, which may introduce variations of up to a few hundred meV 21. We also see the difference between using PseudoDojo and SG15 pseudopotentials in Table 3 and 4. Further differences come from distinct real-space grid treatments and employing the real-time (time-domain) $GW$ builds screening via TDH propagation, in contrast to (truncated) summation over states methods 37. Nevertheless, we can reliably confirm the role of the relativistic effects. The FR–SR correction to the HOMO–LUMO gap is $\Delta E_{g}$ around 0.87 to 0.90 eV regardless of the pseudopotential used (Table 4), and it is consistent with deterministic $G_{0}W_{0}$ benchmarks (0.88 eV) 21.
@@ 4.2.2 Periodic system
For the periodic system, we verify the spin-resolved sGWGW method on bulk AlSb. A dense real-space grid with spacing $dx=0.27$ a.u. is employed, using PseudoDojo 35 PBE SR and FR pseudopotentials. The band gap calculations are shown in Table 5 and are consistent with deterministic $GW$ results 21.
We examine the convergence with respect to $N_{\eta}$, as shown in Figure 2. Both SR and FR pseudopotentials exhibit similar convergence, and $N_{\eta}=8$ is sufficient to obtain an accurate CBM quasiparticle energy $E_{qp}$ as shown in Figure 2 (top). This confirms the stochastic TDH performance analysis: noncollinear nonmagnetic systems require a comparable $N_{\eta}$ to achieve the same error level as collinear systems. The corresponding total core-hour costs are presented in Figure 2 (bottom). In the noncollinear case, the computational effort increases for two reasons: (i) the FFTs must be performed on spinor wavefunctions, effectively doubling their cost, and (ii) the SOC nonlocal pseudopotential terms introduce additional overhead. As a result, the runtime is roughly three times that of the spin-unpolarized case, rather than just twice. Nonetheless, this is still significantly more efficient than deterministic $GW$, where achieving a similar accuracy would require a factor of four increase in additional cost. Further, note that this is a cost increase relative to the spin-degenerate calculation. In the sGWGW calculation, it scales merely linearly with a small prefactor, compared to the conventional calculations. Thus, the reported calculations are orders of magnitude less expensive than those obtained by a conventional route. This suggests that noncollinear sGWGW preserves the favorable linear-scaling behavior observed in spin-unpolarized systems, consistent with the analysis in Refs.16, 18.
@@ 5 Conclusion and outlook
In the future, this unified sGWGW framework will lay the foundation for systematically incorporating spin–spin interactions and vertex corrections. Such extensions will be crucial for accurately capturing excitonic effects, correlated magnetism, and unconventional superconductivity in real materials. Besides, AI-based solvers and learning modules offer a natural path to further accelerate stochastic propagations, particularly by exploiting low-rank structures and transferable features across materials 38, 39, 40, 41. These developments highlight the synergy between stochastic methods and data-driven approaches, opening the door to large-scale quasiparticle studies of large scale complex spin–orbit–coupled and magnetic materials that would otherwise be computationally prohibitive.